### v2

# Spatial, Spectral and Temporal Response of High Intensity Laser Plasma Mirrors–Direct Observation of the Ponderomotive Push

Abstract
Plasma-based optics have emerged as a powerful platform for manipulating and amplifying ultra-intense laser pulses. However, the inherently nonlinear and dynamic nature of plasma leads to significant spatial, spectral, and temporal modulations when driven at relativistic intensities. These modifications can dramatically alter the structure of the reflected laser pulses, posing challenges for their use in applications such as vacuum ultraviolet (VUV) and X-ray generation, as well as relativistic particle acceleration. Comprehensive, multidimensional diagnostics are essential to accurately characterize these so-called ‘plasma mirrors’ (PMs). We present a direct, in situ measurement of the three-dimensional plasma surface evolution during femtosecond laser irradiation, achieved through simultaneous analysis of the wavefront, spectrum, and temporal profile of the reflected light. Our measurements reveal surface deformations on the order of a few hundred nanometers at relativistic intensities, in agreement with three-dimensional particle-in-cell (3D-PIC) simulations. Additionally, the PM induces substantial modifications to the pulse spectrum and temporal profile, introducing spatio-temporal couplings.
I Introduction
Recent advancements in laser technology facilitate the study of the interaction between high-power, ultrahigh contrast femtosecond pulses with solid-density matter [1, 2, 3, 4]. This opens up avenues in high-energy density science, with potential applications in high-brightness vacuum ultraviolet (VUV) and X-ray sources [5], attosecond optics [6], tabletop ultrafast X-ray [7, 8] imaging and ultrafast diffraction [9]. These developments are critically dependent on the temporal and spatial characteristics of femtosecond pulses [10]. Unlike picosecond and nanosecond pulses, where interaction occurs in plasma that expands below solid density, ultrahigh contrast femtosecond pulses induce nearly instantaneous excitation of solid-density matter [11]. In addition, short-pulse, high-intensity lasers can create conditions where quantum electrodynamics (QED) effects [12] and relativistic effects [13] become significant. The extreme pressure (of the order of several 100 Mbar) exerted on the plasma surface leads to surface deformation resulting in various instabilities[14]. Once the plasma is formed, the laser-plasma interaction is predominantly governed by the critical density surface, which acts as a reflective and absorptive interface—referred to as the ‘plasma mirror’ (PM). The dynamic relativistic PM in turn severely modifies the spectral and temporal properties of the laser pulse [6]. Therefore, it becomes important to accurately measure the spatial, spectral and temporal characteristics of these relativistic PMs. At these relativistic intensities, the ponderomotive force (PMF), a nonlinear force arising from the interaction of the laser field and charged particles, plays a crucial role in governing the physical phenomena such as nonlinear self-action effects, particle acceleration, multiphoton ionization and chaotic laser plasma dynamics.
Phenomena such as radiation pressure acceleration (RPA) [15, 16, 17], high harmonic generation (HHG)[6], and laser absorption mechanisms like resonance absorption [18] are highly affected by the curvature introduced by the PMF. Significant progress has been made in probing these ultrafast dynamics. Techniques based on spectral measurements [19, 20, 21, 22, 23] and HHG-based diagnostics [24, 25] have been explored. However, spectrum-based methods are inherently one-dimensional, providing only velocity information, while HHG-based techniques offer two-dimensional insights but fail to directly measure the curved geometry of plasma mirrors. Due to these challenges and the lack of direct measurements, the community often relies on simplified theoretical models and simulations [26, 27, 28, 29, 30, 31]. These approaches, however, are contingent upon various assumptions and initial conditions, which may not fully capture the complexities of the interaction.
Figure 1: a) Schematic of the experimental setup for measuring laser effects on electrons with a QWLSI wavefront sensor as the detector. The figure illustrates the ponderomotive push caused by the intense laser and shows various forces acting on an electron. The PMF on an electron has three components along three directions, as illustrated. $F_{pol}$ is directed along the polarization direction, $F_{k}$ is in the laser propagation direction, and $F_{grad}$ is along the gradient of the average intensity of the field. b) Measured focal spot in vacuum. c) Laser contrast measured using a third-order nonlinear process. d) Diagram of a QWLSI wavefront sensor, which comprises a cross grating and a CCD. e) Example of a raw and zoomed-out portion of an interferogram as measured by the QWLSI wavefront sensor.
In this work, we report the first direct, in situ three-dimensional surface mapping of a plasma mirror during femtosecond pulse interactions by measuring the wavefront of the reflected light from the plasma. The wavefront carries the information of surface depth and provide a direct and easy measure of the plasma surface. The deformation has been measured for different intensities and has been shown to match the theoretical model developed in [24]. We also measured the spectrum and temporal profile of the reflected pulse to gain deeper insight into the underlying plasma dynamics. The reflected pulse undergoes significant modifications in both the spectral and temporal domains. For comparison, an ultrahigh-contrast second harmonic pulse was employed to examine the spatial and spectral properties of the reflected beam. In addition our high-fidelity 3D3V and 2D3V particle-in-cell (PIC) simulations successfully model the laser-induced depression of plasma density surfaces caused by the PMF. Additionally spectral analysis of reflected field indicate the presence of odd higher harmonics.
II Experimental Details
The experiments were performed with the multi-terawatt laser system at Tata Institute of Fundamental Research, Mumbai[32]. A Ti:Sapphire laser (30 fs , 800 nm) with high picosecond contrast ($10^{-9}$) and producing relativistic intensities, was used to create overdense plasma on 5 mm thick solid glass (BK7) targets. The laser was focused with an f/3 off-axis dielectric-coated parabola to a 7 $\mu m$ spot (FWHM) (Figure 1b) to achieve different intensities. The laser reflected from the target plasma was collected and sent outside the vacuum chamber. The beam was imaged on a wavefront measuring device phasics SID-4 (Figure 1a). The beam was also sent to Ocean Optics HR 2000 spectrometer and a single shot temporal profile measuring device GRENOUILLE [33].
We use a high definition wavefront measuring device also known as QuadriWave Lateral Shearing Interferometer (QWLSI), invented by Primot et al. in 2000[34][35]. The QWLSI is composed of a two dimensional transmission grating, a phase checkerboard, and a CCD device. Four replicas of the sample wavefront to be measured are produced and the interference pattern is recorded on a CCD camera. The period of the phase checkerboard is adjusted such that 4 waves are generated from 4 first-order diffraction orders of the incoming wave (see Figure 1d). The wavefront is retrieved from the interferogram by properly masking the spatial frequencies in the Fourier domain and integrating the gradients along the transverse axis. The combination of the grating and phase checkerboard is called a modified Hartmann mask (MHM) due to its similarity with the traditional Hartmann sensors. As QWLSI wavefront devices are very sensitive (5 nm surface RMS) and have a high spatial resolution (27 $\mu m$), these devices are routinely used in adaptive optics setups for wavefront correction in high-intensity lasers where the wavefront becomes heavily distorted by the different amplification stages [36]. They also are used extensively in biology for phase microscopy[37].
Figure 2: a) Wavefront of the reflected laser at different intensities. These are real time 3D surface deformations measured. X and Y represents the spatial dimension and Z is the phase plotted in terms of OPD in vacuum. b) Top left sketch represents two dimensional density profile at low intensity, just start of the pulse, as the intensity increases, the plasma surface gets deformed, piling up of electrons and ions happens, represented by the top right figure. The bottom figure shows the exponential density profile in the x direction. The pulse initially reflects from $n_{c}cos^{2}(\theta)$ where $\theta$ is the incidence angle with respect to the target normal. The cartoon also shows the density bump created by the laser pulse and the total surface deformation. c) Line outs taken at each intensities. Line curves represent the corresponding surface deformation obtained from the theoretical calculation. Shows the experimental data and the model is in good agreement.
The phase obtained from the sensor was converted into the Optical Path Difference (OPD) corresponding to free-space propagation in vacuum, using the relation $OPD=\Phi\lambda/{2\pi}$, where $\Phi$ is the phase and $\lambda$ is the wavelength. Phase measurements were conducted at varying target intensities, creating distinct regimes of laser-matter interaction. At each intensity level, 10 laser shots were taken, and the phases were averaged to reduce numerical error.
III Results and Discussion
It is well known that ponderomotive force (PMF) arises from spatial intensity gradients in a light field [38, 39]. In the non relativistic case, the PMF is a time-averaged force having direction opposite to that of the intensity gradient, and is independent of the particle charge. In general, the PMF $F_{P}$ can be expressed as $F_{P}=-\frac{\partial T_{ik}}{\partial x_{k}}$ where summation is applied on the repeated index and $T_{ik}$ is the Maxwell stress tensor[40]. In the relativistic regime, although the main features of the non-relativistic PMF hold (i.e. electrons expelled form high intensity regions and average motion is independent of the laser polarization), the force mainly depends on the negative gradient of the effective mass[41]. Relativistic treatment has been done theoretically for a focused intense laser[28]. Also calculations have been carried out for electrons in a weakly inhomogeneous light field [42] which typically is the case for a laser focused with a focusing element having focal length much larger than the laser wavelength($f>>\lambda$). In this case the PMF on a electron has three components along three directions as shown in (Figure 1a), $F_{pol}$ is directed along the polarization direction, $F_{k}$ is in the laser propagation direction and $F_{grad}$ is along the gradient of the average intensity of the field.
For a laser with intensity I, the pressure exerted is given by $P=(1+R)I/c$, where R is the reflectivity of the plasma surface. The same relation can be obtained by space integrating the PMF. The expanding plasma surface is pushed into the target when the light pressure exceeds the plasma pressure ($P>nkT_{e}$), which is usually the case for high intensity lasers.
Figure 3: Spectrum of the reflected 800 nm pulse. The corresponding intensities are indicated in units of $W/\mathrm{cm}^{2}$. At higher intensities, the spectrum becomes broadened, and partial absorption is observed near 825 nm.
The proper description of the PMF should include electron and ion motion at the particular intensity. As the electrons respond instantaneously and ions take time to respond due to their heavy mass, the dynamics are seen in two different timescales. The electrons can be assumed to oscillate with the laser field in a constant ion background and having a slow drift corresponding to the ion motion over the full laser duration. The electrons are pulled out and pushed back in every cycle of the laser pulse, forming relativistic electron jets and generating of higher harmonics from the plasma. Although the ion motion in one optical cycle of the laser pulse is negligible, the ions do move significantly over the whole femtosecond laser pulse, which can be seen from the simulations.
Figure 4: Temporal profiles of the reflected fundamental pulse. The color coding follows that of Figure 3.
As seen in Figure 2a, the wavefront becomes increasingly curved, indicating that the plasma surface is pushed towards the target as the intensity increases (note: for better visualization, the plot has been flipped upside down along the Z-axis). The lineouts of the plots are shown in Figure 2c. Due to the PMF, density steepening occurs, which results in the curvature of the plasma surface. After the pulse propagates, the hot plasma begins to expand exponentially toward the vacuum, establishing a density gradient.
A part of the reflected light is also sent to a spectrometer to analyze the spectral behavior of the plasma mirror, as shown in Figure 3. The laser spectrum is significantly modified by the plasma surface. At high intensity, the spectrum broadens, and one part of it (around 825 nm) undergoes heavy absorption. As the intensity decreases, the spectrum starts to resemble the original laser spectrum. The spectral broadening can be explained by self-phase modulation in the plasma leading to possibilities of pulse compression with appropriate chirp compensation. To measure the temporal profile of the reflected laser after the interaction, the collected light was collimated and was sent to a GRENOUILLE (Frequency Resolved Optical Gating) device. Because of the collection optics and chamber window the pulse gets stretched before reaching the detector, but still gives insight about the pulse profile at different intensities.
Figure 5: a) Measured wavefronts of the ultrahigh-contrast pulse. The headings indicate the peak intensities corresponding to each shot, in units of $W/\mathrm{cm}^{2}$. b) Measured focal spot of the 400 nm pulse at the target surface, showing elongation along the $y$-direction. c) Computed temporal contrast of the second harmonic (400 nm) pulse, derived from the measured contrast of the 800 nm fundamental pulse.
Figure 6: Spectrum of the reflected ultrahigh-contrast 400 nm pulse. The spectral bandwidth increases monotonically with intensity. At the highest intensity, the spectrum broadens to nearly three times its initial width, indicating the potential for generating a more compressible pulse.
The experiment was conducted using a high-contrast laser pulse, assuming an exponential density distribution, $n(x)=n_{0}\exp(-x/L)$, where $n_{0}$ represents the solid density, and $L$ is the scale length. The scale length, defined as the position where the density decreases to $1/e$ of the initial value, significantly influences laser absorption, the energy distribution of heated electrons, and high harmonic generation. The scale length $L$ was determined[22] using the frequency domain interferometry technique [43]. Figure 2a presents the measured wavefronts and optical path difference (OPD) maps for various laser intensities, while Figure 2c shows the corresponding line cuts of the OPD maps.
We use the same model developed in [24] to describe our experimental data. The total push by the laser can be described by electron excursion $x_{e}(t)$ from the ion surface, $x_{i}(t)$ which also move significantly over the laser pulse. The total surface deformation is then the given by $x_{T}(t)=x_{i}(t)+x_{e}(t)$. The expressions for the $x_{i}(t)$ and $x_{T}(t)$ are given by,
[…]
x_T(t)=x_i(t)+L ln[1+2 aL(t)(1+sinθ)2 πL / λL ncn0 e^-x_i(t) / L]
where $\Pi_{0}=(R\cos\theta Zm_{e}/2{Am_{p}})^{1/2}$ with $A$ and $Z$ are ion mass number and average charge state, $m_{p}$ and $m_{e}$ are proton mass and electron mass, $\theta$ is the incidence angle, $\lambda_{L}$ is the laser wavelength, and $a_{L}$ corresponds to the laser envelope and is dependent on time and space for a focused laser beam. We fit this model to the experimental data and simulation results, which shows good agreement Figure 2c.
Figure 7: The ion density of the overdense target for non-dimensional laser amplitudes of (a) 2, (b) 10, and (c) 20 is presented after interaction with a relativistic laser. The ponderomotive push results in a cavity whose mean depth increases with amplitude. Additional features such as longitudinal and lateral striations are visible. A comparison of the maximum cavity depth, as measured in the experiment and calculated in the simulation, is shown in (d). The simulation(blue circles) and experimental(purple squares) values show agreement, confirming the logarithmic dependence(simulation fit: dotted blue line).
Figure 8: Zoomed view of (a)$a_{0}=2$ and (b) $a_{0}=4$ and (c,d) cut section views for $a_{0}=2$ of the cavity created by the relativistic pulse are presented.
For comparison, additional experiments were performed using an extreme-contrast second harmonic pulse (400 nm) generated from an LBO crystal [44]. The measured pulse profile and its variation with intensity are detailed in Aparajit et al. [45]. The computed temporal contrast and the experimentally measured focal spot are shown in Figure 5c and Figure 5b, respectively. The measured wavefronts at different intensities are plotted in Figure 5a. The ponderomotive dent is observed to deepen with increasing intensity. The elliptical shape of the wavefront closely reflects the focal spot characteristics of the beam (Figure 5b). The spectrum of the reflected ultrahigh-contrast pulse is presented in Figure 6, showing significant modulation and broadening as the intensity increases. At the highest intensity, the spectrum broadens to nearly three times its initial width, indicating the potential for generating a more compressed pulse.
IV Simulation
To complement our experimental observations, we performed 3D3V and 2D3V Particle-In-Cell (PIC) simulations using the massively parallel, fully relativistic, open-source OSIRIS 4.0 framework [46]. These simulations provide a detailed understanding of plasma dynamics at resolutions and time scales inaccessible in experiments, offering insights into energy deposition, charge transport, and harmonic generation.
The simulations were carried out in a domain of size $300\times 300\times 150$ cells, with a grid resolution of $dx=dy=dz=0.5\omega_{pe}/c$. A linearly-polarized (along $\hat{y}$) laser pulse of wavelength $(\lambda=800)nm$, duration $10fs$, and spot size $1\mu m$ was incident normally ($0^{\circ}$ angle of incidence) on a fully ionized plasma. The laser intensity was varied with normalized field strengths $(a_{0}=0.5,1.0,1.5,2,10,20)$ to study the effects of intensity on plasma evolution ($a_{0}=0.86\lambda_{L}[\mu m]\sqrt{I_{0}[10^{18}W/cm^{2}]}$. for $\lambda_{L}=0.8\mu m,I_{0}[10^{18}W/cm^{2}]=2.14,a_{0}=1.01$ ).
The plasma target was modeled at density $50n_{c}$, consisting of electrons and static neutralizing ions. The ion-to-electron mass ratio was chosen as $m_{i}/m_{e}=25$. The simulation parameters, including laser and plasma properties, are summarized in Table 1 .
Table 1: Simulation Parameters
Parameters Standard Units Normalized Units
Laser
Wavelength $\lambda_{L}$ 800 nm $44[c/\omega_{pe}]$
Frequency $\omega_{L}$ $2.35\times 10^{15}$ rad/s $0.14[\omega_{pe}]$
Intensity $I_{o}$ $5.37\times 10^{17}$ W/cm2 $a_{o}=0.5$
Plasma
Electron Plasma frequency $\omega_{pe}$ $1.66\times 10^{16}$ rad/s $1[\omega_{pe}]$
Critical density $n_{c}(\omega_{L})$ $1.73\times 10^{27}$ 1/m3
Electron number density $n_{e}$ $50n_{c},8.67\times 10^{28}$ 1/m3
Ion number density $n_{i}$ $8.67\times 10^{28}$ 1/m3
Mass of electrons $m_{e}$ $9.109\times 10^{-31}$ kg
Mass of ions $m_{i}$ $25m_{e}$
The depth of the plasma surface depression due to the ponderomotive push exerted by the laser field was analyzed as a function of $a_{0}$, as shown in Figure 7 . The results indicate a monotonic increase in penetration depth with increasing laser intensity, consistent with theoretical predictions of ponderomotive scaling.
The modification of plasma density due to laser interaction is illustrated in Figure 8 , showing snapshots of electron, ion, and charge densities. The charge separation induced by the laser field results in the formation of a characteristic plasma cavity at higher intensities.
To analyze the impact of laser intensity and plasma density on the reflected electromagnetic field, we performed a Fourier analysis of the transverse field component. Notably, the reflected field exhibits odd harmonics, a signature of relativistic plasma boundary oscillations, which arise due to nonlinear plasma interactions[47, 48].
Our PIC simulations validate the experimental findings while providing deeper insights into laser-plasma interactions. The observed trends in ponderomotive push, density modification, and spectral modulation highlight the importance of laser intensity and plasma density in determining the interaction dynamics. These results serve as a benchmark for future high-intensity laser experiments and theoretical investigations.
V Conclusion
We have reported direct measurement of the spatial, spectral, and temporal characteristics of light reflected from relativistic PMs, in agreement with both particle-in-cell (PIC) simulations and theoretical models. We present a direct experimental measurement of the shape of the critical electron density surface in a high intensity laser excited plasma on a solid surface for two different laser wavelengths, providing clear evidence of the ponderomotive push. The behavior of PMs was investigated using high- and ultrahigh-contrast pulses. Significant spectral broadening and pulse shape modification are observed, arising from the deformation of the PM and various nonlinear plasma dynamics. As the conventional metal-coated or dielectric mirrors are unsuitable at relativistic intensities due to their low damage thresholds, relativistic PMs provide a promising alternative for reaching extreme intensities and exploring quantum electrodynamics (QED) regimes. The curvature induced to the reflected beam by the PM can be utilized in the far field for multiple applications, such as controlling the divergence of attosecond pulses generated during the interaction or focusing the reflected beam.